# Experimental Design for Active Transductive Inference in Large Language Models

### v1

@@ Abstract
Transduction, the ability to include query-specific examples in the prompt at inference time, is one of the emergent abilities of large language models (LLMs). In this work, we propose a framework for adaptive prompt design called active transductive inference (ATI). We design the LLM prompt by adaptively choosing few-shot examples for a given inference query. The examples are initially unlabeled and we query the user to label the most informative ones, which maximally reduces the uncertainty in the LLM prediction. We propose two algorithms, GO and SAL, which differ in how the few-shot examples are chosen. We analyze these algorithms in linear models: first GO and then use its equivalence with SAL. We experiment with many different tasks and show that GO and SAL outperform other methods for choosing few-shot examples in the LLM prompt at inference time.
@@ 1 Introduction
Large language models (LLMs), such as Vicuna (Chiang et al., 2023), Falcon-40B (Penedo et al., 2023), and OpenLLaMA (Touvron et al., 2023) are applied in mainly two ways: fine-tuning and prompt tuning. In fine-tuning, the LLM weights are adapted to a downstream task (Devlin et al., 2018). fine-tuning allows for the inclusion of domain-specific knowledge that a pre-trained model may not possess and resembles classic inductive inference. Fine-tuned models often do not need carefully designed prompts, which makes them easier to deploy. The main drawback of fine-tuning is that it can be costly, because tens of thousands of training examples may be needed to fine-tune billions of parameters of the LLM (Ding et al., 2023). In prompt tuning, the LLM weights are fixed and the LLM is given query-specific examples at inference time that affect its output (Lester et al., 2021). This ability to conduct transductive inference is one of the emergent abilities of LLMs. Prompt tuning does not require large training sets. It is also preferred when query-specific examples are private or change over time, and thus can only be utilized at inference time. Prior works on prompt tuning mainly focus on hard prompt tuning, where the prompt is carefully handcrafted to get the desired output. This can be time-consuming, as even a simple change in the prompt can lead to poor performance on the downstream task (Suzgun et al., 2022). We propose a framework for adaptive prompt design called active transductive inference (ATI). The key idea is to design the LLM prompt by adaptively choosing few-shot examples for a given inference query. The examples are initially unlabeled and we query the user to label the most informative ones, which maximally reduce the uncertainty in the LLM prediction. To illustrate our approach, we show two exemplary scenarios next.
The first example is an LLM as a medical diagnostics chatbot. The LLM is tasked with identifying potential medical conditions based on symptoms. The user inputs that they experience fatigue, cough, and shortness of breath. These symptoms can be caused by various conditions: a common cold, flu, or something more severe like heart issues. Instead of immediately offering a range of potential remedies, the LLM could actively query the user with symptoms that are specific to a particular condition, such as chest pain, leg swelling, and irregular pulse for heart-related issues. It then asks the user to label these symptoms as present or absent. This interaction allows the LLM to refine its diagnosis.
@@
Figure 1: A general workflow of active transductive inference (ATI). The active learning module selects examples sequentially to be labeled by a human. Since LLMs can be viewed as general inference machines, our approach is broadly applicable to classification, regression, general pattern recognition, and natural language tasks. The general pattern recognition tasks include Abstract Reasoning Task (ARC) (Alford, 2021) and identifying Probabilistic Context Free Grammar tasks (PCFG) (Hupkes et al., 2020).
To address the above challenges, we propose a framework for adaptive prompt design called active transductive inference (ATI). Our solution is general and can be easily extended to any active supervised-learning task, like active regression (Gao & Koller, 2011) and active classification (Gao & Koller, 2011) (see Figure 1). At a high level, we treat the LLM as a general inference machine that is shown adaptively-chosen examples with labels at inference time. The LLM then utilizes them to answer the original user’s query. The key idea is to choose the next example to label as the one that maximally reduces the estimated uncertainty of the answer to the user’s query. We focus on designing algorithms with the following two properties: (1) Implementable in any LLM that can be queried efficiently. The parameters of the LLM do not change or have to be observed. (2) Analyzable in simple models. In this work, we use linear models to motivate and analyze our algorithms.
@@
(1) We propose a G-Optimal design algorithm (GO). The key idea in GO is to query the user to label the closest examples to the example in the inference task. Our main contribution is the right notion of closeness, based on posterior covariance in a simpler model. GO is implementable with any LLM that can be sampled from, and does not require access to model parameters, feature embeddings of the LLM, or its gradients.
@@
(3) GO is motivated by optimal designs in linear models (Kiefer & Wolfowitz, 1960; Pukelsheim, 2006). This allows us to analyze GO in linear models (Theorem 4). Our proof is a major departure from similar analyses in fixed-budget best-arm identification in bandits (Azizi et al., 2022; Yang & Tan, 2022), for instance because we analyze the discrete allocation problem and each unlabeled point can be labeled at most once. We discuss this in detail right after Theorem 4. SAL is more general than GO because it does not make any linear model assumption in its design. We show that SAL and GO are equivalent in linear models in Theorem 5.
(4) We evaluate GO and SAL on UCI (Markelle Kelly, 1988) and OpenML (Vanschoren et al., 2013) regression and classification tasks, custom NLP datasets, abstract reasoning corpus (ARC) tasks (Alford, 2021; Mirchandani et al., 2023), and Probabilistic Context Free Grammar (PCFG) tasks (Hupkes et al., 2020). GO and SAL consistently outperforms other methods for choosing few-shot examples.
We advance the understanding of active transductive inference in LLMs and develop a practical methodology for adaptive prompt design. This paper is organized as follows. Section 2 introduces the problem setting. Section 3 presents our methods and discusses their properties. Section 4 is devoted to analyzing our methods. Section 5 validates our approach empirically. We review related work in detail in Appendix A. Finally, Section 6 summarizes our contributions and suggests avenues for future work.
@@ 2 Setting
We pose the problem of adaptive prompt design as active learning. We adopt the following standard active learning terminology (Dasgupta, 2005). We have $n$ unlabeled points $\mathcal{X}=\left\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\right\}\subset\mathbb{R}^{d}$. Each point is a vector in $d$ dimensions. The space of labels is $\mathcal{Y}\subseteq\mathbb{R}^{d_{y}}$. Each label is a vector in $d_{y}$ dimensions. Each point $\mathbf{x}\in\mathcal{X}$ has a label $Y\in\mathcal{Y}$. The labels are unknown. The point $\mathbf{x}$ and its label $Y$ are related through an unknown model parameter $\bm{\theta}_{*}$. Specifically, $Y=f(\mathbf{x},\bm{\theta}_{*})+\varepsilon$, where $f$ represents the model and $\varepsilon$ is independent zero-mean noise. The stochasticity of $Y$ allows us to model noisy labels even if $\bm{\theta}_{*}$ was known. One point in $\mathcal{X}$ is designated as a target point $\mathbf{x}_{*}$ and its target label is $Y_{*}=f(\mathbf{x}_{*},\bm{\theta}_{*})+\varepsilon_{*}$. We want to infer it by asking a human to label other points in $\mathcal{X}$.
We illustrate our notation next. In theme recognition, $\mathbf{x}$ is a set of similar items, $Y$ is their theme, and $\bm{\theta}_{*}$ is a model parameter that relates $\mathbf{x}$ and $Y$. Since the model parameter $\bm{\theta}_{*}$ is unknown, the target label $Y_{*}$ is unknown. To reduce its uncertainty, and ultimately learn it, we actively query the user for labels of carefully chosen points $\mathbf{x}\in\mathcal{X}$. We reduce the uncertainty of $Y_{*}$ through that of $\bm{\theta}_{*}$. To understand how we achieve it, we need additional notation. We related the points and their labels, conditioned on previously labeled points, using large language models. Specifically, let $H_{t}=(X_{\ell},Y_{\ell})_{\ell\in[t-1]}$ be the history of $t-1$ previously labeled points, where $X_{\ell}$ is the $\ell$-th labeled point and $Y_{\ell}$ is its label. Let $p(\cdot\mid\mathbf{x},H_{t})$ be the distribution of the label of point $\mathbf{x}$ conditioned on $H_{t}$. Sampling from this distribution is implemented implicitly by the LLM, where $H_{t}$ is provided as $t-1$ few-shot examples in the LLM prompt. Then it is queried for the label of $\mathbf{x}$. We visualize this in Figure 1
A useful way of thinking of $p(\cdot\mid\mathbf{x},H_{t})$ is as a Bayesian posterior distribution (Xie et al., 2021; Deng et al., 2023; Lee et al., 2023) associated with a graphical model $\mathbf{x}\to Y\leftarrow\bm{\theta}_{*}\to H_{t}$. This distribution can be sampled from as follows. First, $\bm{\theta}_{*}$ is sampled conditioned on $H_{t}$, which accounts for model uncertainty. Then $Y$ is sampled conditioned on $\bm{\theta}_{*}$ and $\mathbf{x}$, which accounts for additional label uncertainty. In this model, reducing uncertainty in $\bm{\theta}_{*}$ makes $Y_{*}$ more certain. In LLMs, the model parameter $\bm{\theta}_{*}$ is typically not exposed but sampling $Y\sim p(\cdot\mid\mathbf{x},H_{t})$ can be implemented efficiently. The prior distribution of labels is $p(\cdot\mid\mathbf{x},H_{1})$ where $H_{1}=\left\{\right\}$.
@@
To reduce the uncertainty of $Y_{*}\mid\mathbf{x}_{*},H_{t}$, we need to measure it. One option is $-\mathbb{E}_{Y_{*}\sim p(\cdot\mid\mathbf{x}_{*},H_{t})}[\log p(Y_{*}\mid\mathbf{x}_{*},H_{t})]$. The main issue with entropy is that it is hard to estimate for high-dimensional random variables (Vershynin, 2020), especially without having access to $p(\cdot\mid\mathbf{x}_{*},H_{t})$ beyond sampling from it. Therefore, we opt for $\mathrm{cov}[Y_{*}\mid\mathbf{x}_{*},H_{t}]$, which can be estimated using as few as two samples from $p(\cdot\mid\mathbf{x}_{*},H_{t})$. Specifically, we measure the uncertainty of $Y_{*}\mid\mathbf{x}_{*},H_{t}$ by $\operatorname{tr}(\mathrm{cov}[Y_{*}\mid\mathbf{x}_{*},H_{t}])$ (Pukelsheim, 2006). Note that this is the sum of variances of all entries of $Y_{*}\mid\mathbf{x}_{*},H_{t}$. For scalar $Y_{*}$, this becomes $\mathrm{var}[Y_{*}\mid\mathbf{x}_{*},H_{t}]$, a natural measure of uncertainty in linear regression.
One naive approach to reducing the uncertainty of $Y_{*}$ is to label all points in $\mathcal{X}$ and include them in the LLM prompt. This has two issues. First, it is statistically inefficient because all points in $\mathcal{X}$ must be labeled. Second, since LLM prompts have a limited length, not all labeled points may fit. Therefore, we study how to intelligently choose a small subset of $\mathcal{X}$ of size $T$ that would fit in the prompt.
Our active transductive inference (ATI) problem is formally defined as follows. Fix a budget $T$. We then design a sequential adaptive algorithm over $T$ rounds, where the point $X_{t}$ in round $t\in[T]$ is chosen as a function of history $H_{t}$ up to that round. The label of point $X_{t}$ is $Y_{t}=f(X_{t},\bm{\theta}_{*})+\varepsilon_{t}$, where $\varepsilon_{t}$ is independent zero-mean noise in round $t$, and $Y_{t}$ is initially unknown. This makes our setting strictly more general than choosing the best set of labeled examples for prompt design (Lester et al., 2021; Dong et al., 2022; Zhang et al., 2022a; Liu et al., 2023a; Yu et al., 2023; Liu et al., 2023b; Diao et al., 2023). Our goal is to minimize the uncertainty $\operatorname{tr}(\mathrm{cov}[Y_{*}\mid\mathbf{x}_{*},H_{T+1}])$.
@@ 3 Algorithms
This section has four subsections. In Sections 3.1 and 3.2, we introduce our active learning algorithms. The algorithms are proposed for a single target point with a scalar label. In Section 3.3, we generalize the algorithms to vector labels. In Section 3.4, we extend them to multiple target points.
@@ 3.1 Optimal Design Algorithm
Algorithm 1 G-optimal design (GO)
1:  Input: Target point $\mathbf{x}_{*}$, budget $T$
2:
3:  $\mathcal{L}_{1}\leftarrow\emptyset,\ \mathcal{U}_{1}\leftarrow[n],\ H_{1}\leftarrow()$
4:  for $t=1,\dots,T$ do
5:     $I_{t}\leftarrow\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\mathbf{x}_{*}^{\top}(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}\mathbf{x}_{*}$
6:     $X_{t}\leftarrow\mathbf{x}_{I_{t}}$
7:     Observe label $Y_{t}$ of point $X_{t}$
8:     $\mathcal{L}_{t+1}\leftarrow\mathcal{L}_{t}\cup\{I_{t}\},\ \mathcal{U}_{t+1}\leftarrow\mathcal{U}_{t}\setminus\{I_{t}\}$
9:     $H_{t+1}\leftarrow H_{t}\oplus(X_{t},Y_{t})$
10:  end for
11:
12:  Output: Sample $Y_{*}\sim p(\cdot\mid\mathbf{x}_{*},H_{T+1})$
The key idea is to label similar points $\mathbf{x}$ to the target point $\mathbf{x}_{*}$. The notion of similarity is derived based on insights from a linear model. Specifically, let $Y\sim\mathcal{N}(\mathbf{x}^{\top}\bm{\theta}_{*},\sigma^{2})$ and $\bm{\theta}_{\ast}\sim\mathcal{N}(\bm{\theta}_{0},\mathbf{\Sigma}_{0})$, where $\bm{\theta}_{0}$ and $\mathbf{\Sigma}_{0}$ are the prior mean and covariance of the unknown model parameter $\bm{\theta}_{*}$, respectively. Then for any history $H_{t}$, the posterior variance of $\mathbb{E}[Y_{*}\mid\mathbf{x}_{*},H_{t}]$ is $\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t}\mathbf{x}_{*}$, where
is the posterior covariance of $\bm{\theta}_{*}\mid H_{t}$. This is a classic result for Bayesian inference in linear models (Bishop, 2006).
One approach to minimize the uncertainty of $Y_{*}\mid\mathbf{x}_{*},H_{t}$ is to label points that minimize it greedily. Let $\mathcal{L}_{t}\subseteq[n]$ and $\mathcal{U}_{t}\subseteq[n]$ be the indices of labeled and unlabeled points up to round $t$, respectively. Note that $\mathcal{L}_{t}\cup\mathcal{U}_{t}=[n]$. Then the point that minimizes the uncertainty the most, after it is added to history $H_{t}$, is
$I_{t}=\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\mathbf{x}_{*}^{\top}(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}\mathbf{x}_{*}\,.$ (1)
This algorithm is a greedy solution to the G-optimal optimal design (Pukelsheim, 2006; Katz-Samuels et al., 2021). Hence we call it G-Optimal design and abbreviated as GO. The pseudocode of GO is in Algorithm 1. We use $u\oplus v$ to denote the concatenation of vectors $u$ and $v$. Note that GO does not depend on observed labels $Y_{t}$. Similar optimal designs have been used to select informative samples in active learning (Chaudhuri et al., 2015; Mukherjee et al., 2022), bandits (Fontaine et al., 2021; Mason et al., 2021; Mukherjee et al., 2023a, b, 2024), and reinforcement learning (Wagenmaker et al., 2022).
@@
The time complexity of GO is $O(d^{2}nT)$. This is because the algorithm searches for the best point out of at most $n$ points in $T$ rounds. The evaluation of each point in each round, $\mathbf{x}_{*}^{\top}(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}\mathbf{x}_{*}$, takes $O(d^{2})$ time, because $(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}$ can be implemented in $O(d^{2})$ time using the Sherman-Morrison formula. Note that the LLM is queried only once, in the last step to return $Y_{*}$.

### v2

# Optimal Design for Adaptive In-Context Prompt Tuning in Large Language Models

@@ Abstract
One emergent ability of large language models (LLMs) is that query-specific examples can be included in the prompt at inference time. In this work, we use active learning for adaptive prompt design and call it Active In-context Prompt Design (AIPD). We design the LLM prompt by adaptively choosing few-shot examples from a training set to optimize performance on a test set. The training examples are initially unlabeled and we obtain the label of the most informative ones, which maximally reduces uncertainty in the LLM prediction. We propose two algorithms, GO and SAL, which differ in how the few-shot examples are chosen. We analyze these algorithms in linear models: first GO and then use its equivalence with SAL. We experiment with many different tasks in small, medium-sized, and large language models; and show that GO and SAL outperform other methods for choosing few-shot examples in the LLM prompt at inference time.
@@ 1 Introduction
Large language models (LLMs), such as Vicuna (Chiang et al., 2023), Falcon-40B (Penedo et al., 2023), and OpenLLaMA (Touvron et al., 2023) are applied in mainly two ways: fine-tuning and prompt tuning. In fine-tuning, the LLM weights are adapted to a downstream task (Devlin et al., 2018). Fine-tuning can easily incorporate domain knowledge that a pre-trained model may not possess and resembles classic inductive inference. Fine-tuned models often do not need carefully designed prompts, which makes them easier to deploy. The main drawback of fine-tuning is that it can be costly, because tens of thousands of training examples may be needed to fine-tune billions of parameters of the LLM (Ding et al., 2023). In prompt tuning, the LLM weights are fixed and the LLM is given query-specific examples at inference time that affect its output (Lester et al., 2021). This ability to conduct in-context inference is one of the emergent abilities of LLMs. Prompt tuning does not require large training sets. It is also preferred when query-specific examples are private or change over time, and thus can only be utilized at inference time.
Prior works on prompt tuning mainly focus on hard prompts, which are carefully handcrafted to get the desired output. This can be time-consuming and fragile, as minor prompt modifications can lead to a significant performance drop on the downstream task (Suzgun et al., 2022). In contrast, Zhang et al. (2022a, b) and Diao et al. (2023) explored adaptive prompt design using clustering-based and uncertainty-reducing approaches. While these approaches offer some benefits, we argue that optimal designs (Pukelsheim, 2006; Fedorov, 2013) can outperform them by effectively balancing uncertainty and diversity. Similarly to Zhang et al. (2022a, b) and Diao et al. (2023), we propose a framework for adaptive prompt design called active in-context prompt design (AIPD). The key idea is to design the LLM prompt by adaptively choosing few-shot examples for a set of test examples at inference time. The examples are initially unlabeled and we obtain labels for the most informative ones, which maximally reduce the uncertainty in the LLM prediction for all test examples. We assume that the observed labels are collected from experts (human-in-the-loop) or revealed by an oracle (Dasgupta, 2005; Hanneke et al., 2014). The focus on informativeness and diversity ensures efficient label acquisition by selecting the best examples. This reduces reliance on limited and costly resources such as expert labeling.
@@
(1) We propose a G-Optimal design algorithm (GO). The key idea in GO is to retrieve the examples to label that are closest to the set of test examples in the inference task. Our main contribution is the right notion of closeness, based on posterior covariance in a simpler model. GO is implementable with any LLM that can be sampled from, and does not require access to model parameters, feature embeddings of the LLM, or its gradients.
@@
(3) GO is motivated by optimal designs in linear models (Kiefer and Wolfowitz, 1960; Pukelsheim, 2006). This allows us to analyze GO in linear models (Theorem 1). Our proof is a major departure from similar analyses in fixed-budget best-arm identification in bandits (Azizi et al., 2022; Yang and Tan, 2022), for instance because we directly analyze the discrete allocation problem and each unlabeled example can be labeled at most once. We discuss this in detail right after Theorem 1. SAL is more general than GO because it does not make any linear model assumption in its design. We show that SAL and GO are equivalent in linear models in Theorem 2.
(4) We evaluate GO and SAL on UCI (Markelle Kelly, 1988) and OpenML (Vanschoren et al., 2013) regression and classification tasks, custom NLP datasets, abstract reasoning corpus (ARC) tasks (Alford, 2021; Mirchandani et al., 2023), and Probabilistic Context Free Grammar (PCFG) tasks (Hupkes et al., 2020). GO and SAL consistently outperform other active prompt tuning methods (Zhang et al., 2022a, b; Diao et al., 2023) for choosing few-shot examples in majority of the tasks.
We advance the understanding of active in-context prompt design in LLMs and develop a practical methodology for adaptive prompt design. To our knowledge, this is the first paper that analyzes optimal design based prompting approaches that correctly balance uncertainty and diversity-based sampling as opposed to other existing adaptive prompting-based approaches (Zhang et al., 2022a, b; Diao et al., 2023).
This paper is organized as follows. Section 2 introduces the problem setting. Section 3 presents our methods and discusses their properties. Section 4 is devoted to analyzing our methods. Section 5 validates our approach empirically. We review related work in detail in Appendix A. Finally, Section 6 summarizes our contributions and suggests avenues for future work.
@@ 2 Setting
We pose the problem of adaptive prompt design as active learning. We adopt the following standard active learning terminology (Lewis, 1995; Tong and Koller, 2001; Dasgupta, 2005; Dasgupta et al., 2007; Hanneke et al., 2014). We have a $d$-dimensional feature space $\mathcal{X}\subset\mathbb{R}^{d}$ and a $d_{y}$-dimensional label space $\mathcal{Y}\subseteq\mathbb{R}^{d_{y}}$. A labeled example is a pair $(\mathbf{x},Y)\in\mathcal{X}\times\mathcal{Y}$. The feature vectors and labels are related as $Y=f(\mathbf{x},\bm{\theta}_{*})+\varepsilon$, where $f$ is an underlying model, $\bm{\theta}_{*}$ is its parameter, and $\varepsilon$ is an independent zero-mean noise vector. Our goal is to learn to estimate $f$ on test examples by labeling training examples. We have a budget $T$ on the maximum number of training examples that can be labeled. This constraint can arise due to multiple reasons. For instance, human labels may be necessary and they are naturally costly. Another reason may be that the machine learning model has a limited capacity for including labeled examples, such as the length of prompts in LLMs (Zhang et al., 2022a, b; Diao et al., 2023).
Now we introduce our notation in detail. Denote $[m]=\{1,2,\ldots,m\}$. We have $n$ training examples $\mathcal{X}_{\text{examples}}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}$ and $K$ test examples $\mathcal{X}_{*}=\left\{\mathbf{x}_{*,1},\dots,\mathbf{x}_{*,K}\right\}$. We assume that both sets are related, such as being sampled from the same distribution. The label of the test example $\mathbf{x}_{*,k}$ is $Y_{*,k}$. In our motivating theme recognition example the training examples $\mathbf{x}_{i}$ and test example $\mathbf{x}_{*,k}$ is a concatenation of triplets of movies, and the label $Y_{i}$ or $Y_{*,k}$ is the common theme amongst the triplets respectively. We want to infer $Y_{*,k}$ for all $k\in[K]$ without explicitly modeling the complex function $f$. We model the function using an LLM which we treat as a general inference machine because of its large representation capacity (Brown et al., 2020a; Mirchandani et al., 2023). Specifically, let $H_{t}=\{(X_{\ell},Y_{\ell})\}_{\ell\in[t-1]}$ be a set of $t-1$ previously labeled examples, where $X_{\ell}\in\mathcal{X}_{\text{examples}}$ is the $\ell$-th labeled example and $Y_{\ell}$ is its label. Then we denote by $p(\cdot\mid\mathbf{x},H_{t})$ the distribution over labels of an LLM for a queried example $\mathbf{x}$ when $H_{t}$ is used as few-shot in-context examples. To implement this in the LLM, we simply concatenate $\mathbf{x}$ and $H_{t}$ in context (Zhang et al., 2022a, b; Diao et al., 2023). We know that in-context examples affect the distribution of responses of an LLM (Xie et al., 2021; Suzgun et al., 2022; Deng et al., 2023; Lee et al., 2023). So, the problem of learning $f$ under a budget $T$ can be viewed as selecting $H_{T+1}$ such that $p(Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1})$ is high for all test examples $k\in[K]$. This problem is challenging, especially when the training examples need to be labeled.
@@
To effectively reduce the uncertainty of $Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1}$, we need to quantify it. One possibility it to use the entropy $-\mathbb{E}_{Y_{*,k}\sim p(\cdot\mid\mathbf{x}_{*,k},H_{T+1})}[\log p(Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1})]$. This is problematic because the entropy is hard to estimate for high-dimensional random variables (Vershynin, 2020), especially without having access to $p(\cdot\mid\mathbf{x}_{*,k},H_{T+1})$ beyond sampling from it. This is a shortcoming of recent adaptive prompting techniques (Zhang et al., 2022a, b; Diao et al., 2023). Therefore, we propose using the covariance of $Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1}$ as the uncertainty measure. Specifically, we measure the uncertainty of the $k$-th test example by $\operatorname{tr}(\mathrm{cov}[Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1}])$ and the uncertainty over all test examples by $\max_{k\in[K]}\operatorname{tr}(\mathrm{cov}[Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1}])$. Since the trace of the covariance is the sum of the variances in individual dimensions, our objective can be interpreted as minimizing the maximum variance over the predicted labels of all test examples. This is a natural measure of uncertainty in linear models and corresponding optimal designs (Pukelsheim, 2006; Fedorov, 2013).
Before we present our algorithms, we wanted to outline their general design. Given a budget $T$, we design sequential adaptive algorithms over $T$ rounds, where the example $X_{t}\in\mathcal{X}_{\text{examples}}$ in round $t\in[T]$ is chosen as a function of $H_{t}=\{(X_{\ell},Y_{\ell})\}_{\ell\in[t-1]}$ up to that round. Since $H_{t}$ summarizes past actions of the algorithm, we call it a history. The label of example $X_{t}$ is $Y_{t}=f(X_{t},\bm{\theta}_{*})+\varepsilon_{t}$, where $\varepsilon_{t}$ is an independent zero-mean noise vector in round $t$. Our objective is to minimize the maximum uncertainty over all test examples, $\max_{k\in[K]}\operatorname{tr}(\mathrm{cov}[Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1}])$.
@@ 3 Algorithms
Algorithm 1 G-optimal design (GO)
1:  Input: Training set $\mathcal{X}_{\text{examples}}=\{\mathbf{x}_{i}\}_{i=1}^{n}$, test set $\mathcal{X}_{*}=\{\mathbf{x}_{*,k}\}_{k=1}^{K}$, budget $T$
2:  $\mathcal{L}_{1}\leftarrow\emptyset,\ \mathcal{U}_{1}\leftarrow[n],\ H_{1}\leftarrow\{\}$
3:  for $t=1,\dots,T$ do
4:     $I_{t}=\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\left(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)^{-1}\mathbf{x}_{*,k}\,$
5:     $X_{t}\leftarrow\mathbf{x}_{I_{t}}\in\mathcal{X}_{\text{examples}}$
6:     Observe label $Y_{t}$ of example $X_{t}$
7:     $\mathcal{L}_{t+1}\leftarrow\mathcal{L}_{t}\cup\{I_{t}\},\ \mathcal{U}_{t+1}\leftarrow\mathcal{U}_{t}\setminus\{I_{t}\}$
8:     $H_{t+1}\leftarrow H_{t}\cup\{(X_{t},Y_{t})\}$
9:  end for
10:  Output: Sample $Y_{*,k}\sim p(\cdot\mid\mathbf{x}_{*,k},H_{T+1})$ for $k\in[K]$
In this section, we introduce our active learning algorithms for selecting most informative training examples from $\mathcal{X}_{\text{examples}}$. To simplify notation, we assume scalar labels and then discuss an extension to vector labels at the end of the section. We also let $\mathcal{L}_{t}\subseteq[n]$ and $\mathcal{U}_{t}\subseteq[n]$ be the indices of all labeled and unlabeled training examples up to round $t$, respectively. Note that $\mathcal{L}_{t}\cup\mathcal{U}_{t}=[n]$.
@@ 3.1 Optimal Design Algorithm
The key idea is to label examples in $\mathcal{X}_{\text{examples}}$ that minimize the maximum uncertainty of predictions over all test examples $\mathbf{x}_{*,k}$. Our computation of uncertainty is borrowed from linear models. Specifically, take a linear model $Y=\mathbf{x}^{\top}\bm{\theta}_{*}+\varepsilon$, where $\bm{\theta}_{*}\in\mathbb{R}^{d}$ is its parameter and $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ is independent noise. Suppose that $\bm{\theta}_{\ast}\sim\mathcal{N}(\bm{\theta}_{0},\mathbf{\Sigma}_{0})$. Then a well-known result in Bayesian statistics (Bishop, 2006) is that the posterior variance of the model estimate at an example $\mathbf{x}_{*,k}$ given labeled examples $H_{t}$ is $\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t}\mathbf{x}_{*,k}$, where $\widehat{\mathbf{\Sigma}}_{t}=(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top})^{-1}$ is the posterior covariance of $\bm{\theta}_{*}\mid H_{t}$. Therefore, the maximum uncertainty over test examples is $\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t}\mathbf{x}_{*,k}$. The key observation is that this quantity does not depend on labels. Therefore, it can be optimized greedily by choosing the training example that minimizes it the most,
$I_{t}=\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\left(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)^{-1}\mathbf{x}_{*,k}\,,$ (1)
where $\mathcal{U}_{t}$ are indices of all unlabeled training examples up to round $t$. After the index $I_{t}$ is chosen, the example $\mathbf{x}_{I_{t}}$ and its label $Y_{t}$ are added to the history to get $H_{t+1}$ for the next iteration $t+1$.
This algorithm is a greedy solution to the G-optimal design (Pukelsheim, 2006; Katz-Samuels et al., 2021). We call it G-Optimal design and abbreviate it as GO. The pseudocode of GO is in Algorithm 1. Note that GO does not depend on observed $Y_{t}$. Similar optimal designs have been effectively applied in active learning (Chaudhuri et al., 2015; Mukherjee et al., 2022), bandits (Fontaine et al., 2021; Mason et al., 2021), and reinforcement learning (Wagenmaker et al., 2022). However, this is the first paper that studies optimal design for adaptively designing prompts (Zhang et al., 2022a; Diao et al., 2023). (1) can be viewed as choosing that training example $\mathbf{x}_{i}\in\mathcal{U}_{t}$ that minimizes the maximum eigenvalue of the posterior covariance $\widehat{\mathbf{\Sigma}}_{t}$. Therefore this leads to reducing the uncertainty over the model parameter $\bm{\theta}_{*}$ as the confidence ellipsoid around $\bm{\theta}_{\star}$ shrinks (Lattimore and Szepesvári, 2020). Note that maximum eigenvalue reduction also ensures diversity as it leads to choosing training examples along all directions in $\mathbb{R}^{d}$.
@@
The time complexity of GO is $O(Kd^{2}nT)$. This is because, for $T$ rounds, the algorithm searches for the best training example out of at most $n$ and evaluates it on all test examples $\mathbf{x}_{*,k}\in\mathcal{X}_{*}$. The evaluation of each test example in round $t$, $\mathbf{x}_{*,k}^{\top}(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}\mathbf{x}_{*,k}$, takes $O(d^{2})$ time, because $(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}$ can be computed in $O(d^{2})$ time using the Sherman-Morrison formula. In the last step, the LLM is queried $K$ times to return $\{Y_{*,k}\}_{k=1}^{K}$.